If the square of a positive integer is added to 3 times the integer, the results is 40. Find the integer.

Answer by mananth(16949)

let the number be x
x^2+3x=40
x^2+3x-40=0
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x+8)(x-5)=0
x=-8 OR x=5
but number is positive
so the number is 5